# WebDevJudge: Evaluating (M)LLMs as Critiques for Web Development Quality

### v1

@@ B.3 Agentic Workflow Implementation
Error Type Proportion Example Description Error Reason Non-funtional element 35.9 Task: On a Question Review page, modify the text content of a displayed question within its editable text field. Unable to edit question content. Expected Result: The displayed question text updates to reflect the new input. Missing element 30.9 Task: Click the ‘Market’ tab or navigation item. Unable to find the Market interface. Expected Result: The ‘Market’ view is displayed, showing a list of cryptocurrencies. Prerequisite not met 15.2 Task: Select an answer for a multiple-choice question by clicking on one of the options. Unable to import the quiz correctly. Expected Result: The chosen answer option is visually highlighted. Loading issue 9.0 Task: Click a file name in the File Explorer sidebar. Unable to load file content. Expected Result: The corresponding file’s mock content is shown in an editor tab. Unreasonable outcome 5.4 Task: Scroll the chat display area to view older messages when content overflows. N/A Expected Result: The user is logged out. Ambiguous input 2.2 Task: Hover over a data point or segment in a Chart. Chart is not required in the webpage. Expected Result: A tooltip containing mock data relevant to the hovered chart element is displayed. Overly detailed task 0.9 Task: From a list of activities on the ‘Activities Screen’, click on a specific activity’s name or image. Unable to get the expected result based on the operation. Expected Result: The ‘Detail View’ for the selected activity is displayed, showing its full description. Missing animation 0.5 Task: Move the mouse cursor across the main 3D animated background area. The background 3D animation is not responsive. Expected Result: The 3D background animation dynamically responds to the cursor’s position or movement.
Appendix C WebDevJudge-Unit Dataset
C.1 Dataset Construction
C.2 Dataset Statistics and Examples
The resulting dataset comprises 502 tasks derived from 105 unique queries. The feasibility labels are distributed as follows: 279 tasks are marked as feasible, and 223 as infeasible. Table 10 presents a detailed breakdown of the error types for the infeasible tasks, along with illustrative examples for each category.
Appendix D Additional Results
@@
D.1 Prediction Distribution Consistency
Furthermore, the high consistency in pairwise comparison suggests that while different models may lack a shared understanding of ’absolute quality,’ their internal mechanisms for making ‘relative preference’ judgments are more aligned. This reinforces our view that leveraging relative judgments is a more robust approach for automated evaluation in domains characterized by nuance and ambiguity, where clear-cut right or wrong answers are scarce.
D.2 Mitigating the Impact of Positional Bias
@@
Claude-4-sonnet 65.0 65.1 66.1 66.1
GPT-4.1 64.8 65.1 64.2 66.1
DeepSeekV3-0324 60.8 62.2 62.5 63.6
@@
D.3 Evaluation Results on Non-Tie Cases
@@
Model/Method Digital Design Game & App Web & Special Average Single Pair Single Pair Single Pair Single Pair Likert  GPT-4.1 65.73 80.34 71.27 82.87 72.61 85.35 69.77 82.75  GPT-4o 58.99 78.65 64.09 80.11 70.70 80.89 64.34 79.84  Qwen-2.5-VL-72B-Inst. 50.00 74.72 51.93 76.24 59.87 78.98 53.68 76.55  Gemini-2.5-flash-lite 56.74 69.10 56.91 69.61 57.32 73.89 56.98 70.74 DeepSeek-V3-0324 67.42 78.09 64.09 75.14 71.34 83.44 67.44 78.68 Kimi-K2-Inst. 69.66 78.65 66.30 73.48 63.69 80.25 66.67 77.33 Qwen3-235B-A22B-Inst. 62.36 75.28 67.40 77.35 67.52 75.80 65.70 76.16 Qwen3-30B-A3B-Inst. 65.73 76.97 69.06 73.48 67.52 81.53 67.44 77.13 Qwen2.5-32B-Inst. 55.06 71.91 58.56 77.35 59.24 77.71 57.56 75.58 Gemma-3-27B-it 51.12 69.66 57.46 64.09 49.04 75.16 52.71 69.38 Qwen2.5-14B-Inst. 51.69 59.55 50.28 56.35 43.31 59.87 48.64 58.53  Claude-4-Sonnet 61.24 80.90 69.06 82.32 69.43 80.25 66.47 81.20  Claude-3.7-Sonnet 75.28 79.78 69.61 75.69 69.43 85.35 71.51 80.04  Gemini-2.5-pro 59.55 75.28 58.56 79.01 50.32 73.25 56.40 75.97 GLM-4.5 67.42 79.78 63.54 76.80 64.33 78.98 65.12 78.49 DeepSeek-R1-0528 57.87 76.97 65.75 76.80 58.60 76.43 60.85 76.74 Qwen3-30B-A3B-Think. 62.36 71.35 61.88 72.38 61.78 75.80 62.02 73.06 Rubric  GPT-4.1 78.09 82.02 70.72 73.48 70.70 87.26 73.26 80.62  GPT-4o 75.84 78.09 71.27 75.69 73.25 83.44 73.45 78.88  Qwen-2.5-VL-72B-Inst. 70.79 76.40 71.27 77.9 73.25 80.89 71.71 78.29 DeepSeek-V3-0324 78.09 79.21 65.75 79.01 78.34 87.26 73.84 81.59 Kimi-K2-Inst. 72.47 82.02 67.40 74.59 75.16 87.26 71.51 81.01 Qwen3-235B-A22B-Inst. 74.16 76.97 65.75 66.3 71.97 75.16 70.54 72.67 DeepSeek-R1-0528 78.09 79.21 64.09 70.17 73.25 77.07 71.71 75.39 GLM-4.5 76.97 80.34 66.30 70.72 77.71 79.62 73.45 76.74 Agentic workflow 67.98 - 71.27 - 75.80 - 71.51 -
Appendix E The use of large language models

### v2

@@ C.3 Agentic Workflow Implementation
C.4 Prompt Templates
The prompt templates used for pairwise comparison experiments are as follows. For single-answer grading, the only action is simply to modify the input presentation, for example, by modifying two separate code blocks into one.
Prompt for direct comparison You are tasked with comparing two React code snippets (Model A and Model B) based on the user’s query. The input will contain: - User Query: The question or concern the user has. - Answer from Model A: The output of Model A. - Answer from Model B: The output of Model B. Your output should be in the following JSON format: ‘‘‘ { “reason”: “Detailed explanation for why one model is better, based on the user’s query”, “winner”: “‘model_a’, ‘model_b’, or ‘tie”’ } ‘‘‘ - reason: A detailed explanation of why one model’s answer is better than the other, based on the user’s query. Focus on the quality of the responses and how well they address the user’s concerns.
- winner: Select the winner from the following options:
- “model_a”: If Model A is better.
- “model_b”: If Model B is better.
- “tie”: If both models are equally good or both provide unsatisfactory answers.
Ensure that you thoroughly evaluate the responses before selecting the winner and providing the reason.
Prompt for rubric-based comparison You are an expert Quality Assurance engineer specializing in web development. Your objective is to meticulously evaluate and compare two different web development solutions for the same task based on a predefined rubric. You will be provided with the user’s initial query, two solutions (codes for both webpages A and B), and a comprehensive rubric covering intention, static, and dynamic elements of the webpage. Based on these inputs, you will assess whether each requirement in the rubric is implemented in each of the two solutions. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Be as objective as possible. During your assessment, please note that the solution might use different terminology than the rubric. Consider a requirement met if the solution’s feature is equivalent. For example, the required heading element is present on the webpage, though the exact text or symbol differs. ## User Query {user_query} ## Code A ‘‘‘tsx
{code_a}
## Code B ‘‘‘tsx
{code_b}
## Rubric ### Intention ‘‘‘json
{intention_rubric} ‘‘‘
### Static Elements ‘‘‘json
{static_rubric}
### Dynamic Elements ‘‘‘json
{dynamic_rubric}
## INSTRUCTIONS Your task is to return a single JSON object. This object should have three top-level keys: “intention”, “static”, and “dynamic”. The value for each key should be a JSON object that mirrors the structure of the corresponding rubric provided above. For each leaf node in each rubric (i.e., where “children” is null), you must add a new key ”value”. The value for this key must be a string: “A” if solution A is better, “B” if solution B is better, or “tie” if they are of equal quality or both fail to meet the requirement. ## Output Format Begin your evaluation by providing an explanation for your reasoning. End your output with a JSON object wrapped with ‘‘‘json at the beginning and ‘‘‘ at the end. Do not include any other text after the JSON object. Here is an example of the output format: ‘‘‘json
{ ”intention”: { ”description”: ”The purpose of the web page.”, ”children”: [ { ”description”: ”A web page for book reviews.”, ”children”: null, ”value”: ”A” } ] }, ”static”: { ”description”: ”The static elements of the web page.”, ”children”: [ { ”description”: ”The book review submission form.”, ”children”: [ { ”description”: ”A field to input the book’s rating.”, ”children”: null, ”value”: ”tie” } ] } ] }, ”dynamic”: { ”description”: ”The interaction between the user and the web page.”, ”children”: [ { ”description”: ”Basic user interactions.”, ”children”: [ { ”description”: ”User can type text into the review text area.”, ”children”: null, ”value”: ”B” } ] } ] } } “‘ ”””
Prompt for Likert scale-based comparison You are an expert Quality Assurance engineer specializing in web development. Your objective is to meticulously evaluate and compare two different web development solutions for the same task. You will be provided with the user’s initial query and the solutions (codes for both webpages A and B). Based on these inputs, you will assess the quality of each solution across several key dimensions. For each sub-criteria, you must provide a rating on a 5-point Likert scale, where 1 represents “Very Poor” and 5 represents “Excellent”. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Be as objective as possible. {DIMENSIONS OF THE LIKERT SCALE}
## User Query {user_query} ## Code A ‘‘‘tsx
{code_a}
## Code B ‘‘‘tsx
{code_b} ‘‘‘
## Output Format Begin your evaluation by providing a short explanation. End your output with a json object wrapped with ‘‘‘json at the beginning and ‘‘‘ at the end. Use the sub-criterion id as the key. The value for each key should be a nested json object containing the scores for each solution, with “A” and “B” as keys. Do not include any other text after the json object. Here is an example of the output format: ‘‘‘json
{ ”1.1”: { ”A”: 5, ”B”: 4 }, ”1.2”: { ”A”: 4, ”B”: 5 }, ”1.3”: { ”A”: 3, ”B”: 3 }, ”1.4”: { ”A”: 2, ”B”: 2 }, ”2.1”: { ”A”: 5, ”B”: 5 }, ”2.2”: { ”A”: 4, ”B”: 3 }, ”2.3”: { ”A”: 3, ”B”: 4 }, ”3.1”: { ”A”: 5, ”B”: 5 }, ”3.2”: { ”A”: 4, ”B”: 4 }, ”3.3”: { ”A”: 3, ”B”: 3 }, ”4.1”: { ”A”: 5, ”B”: 4 }, ”4.2”: { ”A”: 4, ”B”: 5 }, ”4.3”: { ”A”: 3, ”B”: 4 } }
Appendix D WebDevJudge-Unit Dataset
D.1 Dataset Construction
@@
Error Type Proportion Example Description Error Reason Non-funtional element 35.9 Task: On a Question Review page, modify the text content of a displayed question within its editable text field. Unable to edit question content. Expected Result: The displayed question text updates to reflect the new input. Missing element 30.9 Task: Click the ‘Market’ tab. Can not find the Market tab. Expected Result: The ‘Market’ view is displayed, showing a list of cryptocurrencies. Prerequisite not met 15.2 Task: Select an answer for a multiple-choice question by clicking on one of the options. Unable to import the quiz correctly. Expected Result: The chosen answer option is visually highlighted. Loading issue 9.0 Task: Click a file name in the File Explorer sidebar. Unable to load file content. Expected Result: The corresponding file’s mock content is shown in an editor tab. Unreasonable outcome 5.4 Task: Scroll the chat display area to view older messages when content overflows. N/A Expected Result: The user is logged out. Ambiguous input 2.2 Task: Hover over a data point or segment in a Chart. Chart is not required in the webpage. Expected Result: A tooltip containing mock data relevant to the hovered chart element is displayed. Overly detailed task 0.9 Task: From a list of activities on the ‘Activities Screen’, click on a specific activity’s name or image. Unable to get the expected result based on the operation. Expected Result: The ‘Detail View’ for the selected activity is displayed, showing its full description. Missing animation 0.5 Task: Move the mouse cursor across the main 3D animated background area. The background 3D animation is not responsive. Expected Result: The 3D background animation dynamically responds to the cursor’s position or movement.
D.2 Dataset Statistics and Examples
The resulting dataset comprises 502 tasks derived from 105 unique queries. The feasibility labels are distributed as follows: 279 tasks are marked as feasible and 223 as infeasible. Table 12 presents a detailed breakdown of the error types for the infeasible tasks, along with illustrative examples for each category.
Appendix E Additional Results
@@
E.1 Mitigating the Impact of Positional Bias
@@
Claude-4-sonnet 68.6 68.3 70.0 70.2
GPT-4.1 68.8 68.3 68.7 70.3
DeepSeekV3-0324 64.2 64.4 66.7 67.1
@@
E.2 Prediction Distribution Consistency
Furthermore, the high consistency in pairwise comparison suggests that while different models may lack a shared understanding of ‘absolute quality,’ their internal mechanisms for making ‘relative preference’ judgments are more aligned. This reinforces our view that leveraging relative judgments is a more robust approach for automated evaluation in domains characterized by nuance and ambiguity, where clear-cut right or wrong answers are scarce.
To further substantiate these findings, we extend our consistency analysis to rubric-based grading, examining its internal consistency across different models (Figure 7) and its alignment with the direct evaluation paradigm (Figure 8).
(a) Consistency under pairwise comparison.
(b) Consistency under single-answer comparison.
Figure 7: The inner prediction consistency between model predictions under rubric paradigm.
(a) Prediction consistency under rubric (pairwise comparison) and direct paradigm.
(b) Prediction consistency under rubric(single-answer grading) and direct paradigm.
Figure 8: The inter-method consistency between model predictions under rubric paradigm and direct paradigm.
E.3 Evaluation Results on Non-Tie Cases
@@
Model/Method Digital Design Game & App Web & Special Average Single Pair Single Pair Single Pair Single Pair Likert  GPT-4.1 67.39 83.7 72.16 84.02 69.46 82.63 69.72 83.49  GPT-4o 58.7 80.98 64.95 81.96 68.26 77.84 63.85 80.37  Qwen-2.5-VL-72B-Inst. 49.46 76.09 53.61 76.29 57.49 76.65 53.39 76.33  Gemini-2.5-flash-lite 57.07 70.65 57.73 70.62 55.69 70.66 56.88 70.64 DeepSeek-V3-0324 67.93 79.35 65.98 76.8 68.86 80.84 67.52 78.9 Kimi-K2-Inst. 69.02 80.98 69.07 76.29 62.87 77.84 67.16 78.35 Qwen3-235B-A22B-Inst. 63.59 77.17 69.07 78.87 66.47 73.65 66.42 76.7 Qwen3-30B-A3B-Inst. 67.93 79.35 70.62 74.74 66.47 79.04 68.44 77.61 Qwen2.5-32B-Inst. 55.43 73.37 56.7 76.8 57.49 75.45 56.51 75.23 Gemma-3-27B-it 50.54 70.65 58.25 64.95 48.5 73.05 52.66 69.36 Qwen2.5-14B-Inst. 51.63 59.78 51.03 57.73 42.51 58.08 48.62 58.53  Claude-4-Sonnet 63.04 84.24 70.1 82.47 69.46 79.64 67.52 82.2  Claude-3.7-Sonnet 76.09 82.07 71.65 76.8 68.26 83.83 72.11 80.73  Gemini-2.5-pro 59.78 76.63 59.28 80.41 51.5 72.46 57.06 76.7 GLM-4.5 67.39 82.61 65.46 78.87 62.28 77.84 65.14 79.82 DeepSeek-R1-0528 59.24 80.98 67.53 79.38 57.49 73.65 61.65 78.17 Qwen3-30B-A3B-Think. 60.33 74.46 61.86 73.71 60.48 74.85 60.92 74.31 Rubric  GPT-4.1 79.89 81.52 72.16 75.77 71.26 87.43 74.5 81.28  GPT-4o 79.35 80.43 72.68 77.32 72.46 83.23 74.86 80.18  Qwen-2.5-VL-72B-Inst. 72.28 78.8 74.23 80.41 73.65 79.64 73.39 79.63 DeepSeek-V3-0324 79.89 83.15 68.56 80.93 77.25 86.23 75.05 83.3 Kimi-K2-Inst. 74.46 83.15 70.1 78.35 75.45 83.83 73.21 81.65 Qwen3-235B-A22B-Inst. 78.8 78.26 69.59 69.07 71.86 74.85 73.39 73.94 DeepSeek-R1-0528 78.8 81.52 67.53 72.16 73.65 77.25 73.21 76.88 GLM-4.5 79.35 82.07 69.07 73.71 77.25 79.64 75.05 78.35 Agentic workflow 70.11 - 67.01 - 76.05 - 70.83 -
Appendix F Case Study
F.1 Examples of Rubric Trees
Examples of the rubric tree structure are illustrated in Figure 9. We can observe that for queries requiring specific domain knowledge (e.g., the mid one), large language models can generate detailed and accurate rubrics, such as evaluation criteria for specific components. This highlights a key advantage of LLMs over humans: the ability to leverage their extensive domain knowledge to generate precise rubrics for corresponding queries. However, for more general queries (e.g., the rightmost query), the rubrics generated by LLMs can be overly fine-grained. This may introduce ambiguity during evaluation, for instance, by adding overly specific categories like “Content” for a general “categories”.
Figure 9: Examples of the rubric tree structure derived from different sources with corresponding queries. The diagram illustrates human-written rubric trees used for few-shot generation (left), alongside good (center) and suboptimal (right) examples of LLM-generated rubric trees. Note that the children attribute of leaf nodes is explicitly set to null. For conciseness, some internal nodes have been omitted. Leaf nodes are labeled in green.
F.2 Failure Cases of LLM-based and Agentic Evaluators
In this section, we present failure cases that illustrate the different failure modes discussed in Section 4.
Figure 10: Failure cases in functional equivalence identification. While GPT-4.1 correctly identifies equivalence for the specific “Organization” category and the general “categories” requirement given the query, rubric, and instructions, other models like DeepSeek-R1-0528 and UI-TARS-1.5 fail to do so. These models misinterpret the requirements and exhibit hesitation in recognizing the equivalence for general categories.
Figure 11: Failure case in feasibility analysis due to operation error. The agentic evaluator fails to detect the target ‘regenerate’ element in the provided screenshot, resulting in an incorrect assessment of the task as infeasible.
Regarding limitations in understanding functional equivalence, we examine the query and rubric shown in Figure 9 (right). Although the rubric presents some ambiguity, human annotators and GPT-4.1 successfully identify functional equivalence for both the specific category “Organization” and the general “categories” requirement, provided with the query, rubric, and instructions. In contrast, other models, such as DeepSeek-R1-0528 and UI-TARS-1.5, struggle with interpretation. As shown in Figure 10, these models suffer from misinterpretation, which leads to incorrect functional equivalence identification.
Regarding limitations in feasibility analysis, agentic evaluators are constrained by their inherent capabilities. Figure 11 illustrates such a case where the agent fails to locate the target ‘regenerate’ element within the screenshot, incorrectly leading it to classify the task as infeasible.
Appendix G The use of large language models